OpenStudy (anonymous):

solve the triangle. B = 36°, a = 41, c = 17

OpenStudy (campbell_st):

well you can draw a diagram  now it depends on what you need to find...? and angle A or C or side b

OpenStudy (anonymous):

can you walk me through this please

OpenStudy (anonymous):

im suffering

OpenStudy (campbell_st):

ok... so lets start by finding b... you need to rule of cosines \[b^2 = 17^2 + 41^2 - 2 \times 17 \times 41 \times \cos(36)\]

OpenStudy (anonymous):

okay i'm tryiing to work it out hold on

OpenStudy (anonymous):

i got 289+1681x1394 cos 36

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

b=465?

OpenStudy (anonymous):

i mean 21.5

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well I got 29.02 I've checked it and think I'm right... so next to find an angle... say angle A use the law of sines \[\frac{\sin(A)}{41} = \frac{\sin(36)}{29.02}\] now solve for A